OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y = x – 2 that passes through the point (4,3)? y = –x – 2 y = x + 6 y = –x + 12 y = –x – 12

OpenStudy (ajprincess):

The gradient of the required line*The gradient of the given line=-1 That is m*m1=-1 m1=1 m*1=-1 m=-1 \[m=\frac{ y-y_1 }{ x-x_1 }\] \[-1=\frac{ y-3 }{ x-4}\] Can u find nw?